# Subsampled Ensemble Can Improve Generalization Tail Exponentially

### v3

@@ Appendix D Supplementary material for numerical experiments
This section supplements Section 3. We first provide details for the architecture of the neural networks in Section D.1, and the considered stochastic programs in Section D.2. Section D.3 presents a comprehensive profiling of hyperparameters of our methods, and Section D.4 provides additional experimental results.
D.1 MLP architecture
@@
D.2 Stochastic programming problems
Resource allocation (Kleywegt et al., 2002)   The decision maker wants to choose a subset of $m$ projects. A quantity $q$ of low-cost resource is available to be allocated, and any additional resource can be obtained at an incremental unit cost $c$. Each project $i$ has an expected reward $r_{i}$. The amount of resource required by each project $i$ is a random variable, denoted by $W_{i}$. We can formulate the problem as
@@ D.2 Stochastic programming problems
Supply chain network design (Shapiro et al., 2021, Chapter 1.5)   Consider a network of suppliers, processing facilities, and customers, where the goal is to optimize the overall supply chain efficiency. The supply chain design problem can be formulated as a two-stage stochastic optimization problem
@@
where $N,C,S$ are appropriate matrices that describe the network flow constraints. More details about this example can be found in (Shapiro et al., 2021, Chapter 1.5). In our experiment, we consider the scenario of 3 suppliers, 2 facilities, 3 consumers, and 5 products. We choose supply $s$ and demand $d$ as random variables that follow the Pareto distribution.
Maximum weight matching and stochastic linear program    We explore both the maximum weight matching problem and the linear program that arises from it. Let $G=(V,E)$ be a general graph, where each edge $e\in E$ is associated with a (possibly) random weight $w_{e}$. For each node $v\in V$, denote $E(v)$ as the set of edges incident to $v$. Based on this setup, we consider the following linear program
@@
Mean-variance portfolio optimization    Consider constructing a portfolio based on $m$ assets. Each asset $i$ has a rate of return $r_{i}$ that is random with mean $\mu_{i}$. The goal is to minimize the variance of the portfolio while ensuring that the expected rate of return surpasses a target level $b$. The problem can be formulated as
$\begin{array}[]{ll}\min_{\theta}&\mathbb{E}\left[(\sum_{i=1}^{m}(r_{i}-\mu_{i})\theta_{i})^{2}\right]\\[5.0pt] \text{subject to }&\sum_{i=1}^{m}\mu_{i}\theta_{i}\geq b,\\[5.0pt] &\sum_{i=1}^{m}\theta_{i}=1,\\[5.0pt] &\theta_{i}\geq 0\ \forall i=1,\ldots,m\end{array}$ (55)
@@ D.3 Hyperparameter profiling
Subsample size    We explored scenarios where $k$ (equivalently $k_{1}$ and $k_{2}$) is both dependent on and independent of the total sample size $n$ (see Figures 5(a), 6(a), and 6(b)). The results suggest that a constant $k$ generally suffices, although the optimal $k$ varies by problem instance. For example, Figures 6(a) and 6(b) show that $k=2$ yields the best performance; increasing $k$ degrades results. Conversely, in Figure 5(a), $k=2$ proves inadequate, with larger $k$ delivering good results. The underlying reason is that the effective performance of $\mathsf{MoVE}$ requires $\theta^{*}\in\operatorname*{arg\,max}_{\theta\in\Theta}p_{k}(\theta)$. In the former, this is achieved with only two samples, enabling $\mathsf{MoVE}$ to identify $\theta^{*}$ with a subsample size of $2$. For the latter, a higher number of samples is required to meet this condition, explaining the suboptimal performance at $k=2$. In Figure 7, we simulate $p_{k}(\theta)$ for the two cases, which further explains the influence of the subsample size.
Ensemble size   In Figure 8, we illustrate the performance of $\mathsf{MoVE}$ and $\mathsf{ROVE}$ under different $B,B_{1},B_{2}$, where we set $k=k_{1}=k_{2}=10$ and $\epsilon=0.005$. From the figure, we find that the performance of our ensemble methods is improving in the ensemble sizes.
Threshold $\epsilon$   The optimal choice of $\epsilon$ in $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ is problem-dependent and related to the number of (near) optimal solutions. This dependence is illustrated by the performance of $\mathsf{ROVE}$ shown in Figures 5(b) and 5(c). Hence, we propose an adaptive strategy defined as follows: Let $g(\epsilon):=1/B_{2}\cdot\sum_{b=1}^{B_{2}}\mathbbm{1}(\hat{\theta}_{n}(\epsilon)\in\widehat{\Theta}^{\epsilon,b}_{k_{2}})$, where we use $\hat{\theta}_{n}(\epsilon)$ to emphasize the dependency of $\hat{\theta}_{n}$ on $\epsilon$. Then, we select $\epsilon^{*}:=\min\left\{\epsilon:g(\epsilon)\geq 1/2\right\}$. By definition, $g(\epsilon)$ is the proportion of times that $\hat{\theta}_{n}(\epsilon)$ is included in the “near optimum set” $\widehat{\Theta}^{\epsilon,b}_{k_{2}}$. The choice of $\epsilon^{*}$ makes it more likely for the true optimal solution to be included in the “near optimum set”, instead of being ruled out by suboptimal solutions. Practically, $\epsilon^{*}$ can be efficiently determined using a binary search as an intermediate step between Phases I and II. To prevent data leakage, we compute $\epsilon^{*}$ using $\mathbf{z}_{1:\lfloor\frac{n}{2}\rfloor}$ (Phase I data) for $\mathsf{ROVEs}$. From Figure 5, we observe that this adaptive strategy exhibits decent performance for all scenarios. Similar behaviors can also be observed for $\mathsf{ROVEs}$ in Figure 9.
D.4 Additional experimental results
@@
Figure 17: Comparison of $\bar{\eta}_{k,\delta}$ for $\mathsf{MoVE}$ and $\mathsf{ROVE}$ in a linear program with multiple optima (corresponds to the instance in Figure 3(e)). Threshold $\epsilon$ is chosen as $\epsilon=4$ when $k=k_{1}=k_{2}=10$ and $\epsilon=2.5$ when $k=k_{1}=k_{2}=50$, according to the adaptive strategy. Note that $\bar{\eta}_{k,\delta}=\max_{\theta\in\Theta}p_{k}(\theta)-\max_{\theta\in\Theta\backslash\Theta^{\delta}}p_{k}(\theta)$ by (29), which measures the generalization sensitivity. For $\mathsf{MoVE}$, we have $p_{k}(\theta)=\mathbb{P}(\hat{\theta}_{k}^{SAA}=\theta)$; and for $\mathsf{ROVE}$, we have $p_{k}(\theta)=\mathbb{P}(\theta\in\widehat{\Theta}_{k}^{\epsilon})$, where $\widehat{\Theta}_{k}^{\epsilon}$ is the $\epsilon$-optimal set of SAA defined in (28). From the figure, we can observe that the issue brought by the presence of multiple optimal solutions can be alleviated using the two-phase strategy in $\mathsf{ROVE}$.

### v4

@@ Appendix F Supplementary Material for Numerical Experiments
This section supplements Section 3. We first provide details for the architecture of the neural networks in Section F.1, and the considered stochastic programs in Section F.2. Section F.3 presents a comprehensive profiling of hyperparameters of our methods, and Section F.4 provides additional experimental results that evaluate our algorithms from various perspectives.
F.1 MLP Architecture
@@
F.2 Stochastic Programming Problems
Resource Allocation (Kleywegt et al., 2002).
The decision maker wants to choose a subset of $m$ projects. A quantity $q$ of low-cost resource is available to be allocated, and any additional resource can be obtained at an incremental unit cost $c$. Each project $i$ has an expected reward $r_{i}$. The amount of resource required by each project $i$ is a random variable, denoted by $W_{i}$. We can formulate the problem as
@@ F.2 Stochastic Programming Problems
Supply chain network design (Shapiro et al., 2021, Chapter 1.5).
Consider a network of suppliers, processing facilities, and customers, where the goal is to optimize the overall supply chain efficiency. The supply chain design problem can be formulated as a two-stage stochastic optimization problem
@@
where $N,C,S$ are appropriate matrices that describe the network flow constraints. More details about this example can be found in Shapiro et al. (2021, Chapter 1.5). In our experiment, we consider the scenario of 3 suppliers, 2 facilities, 3 consumers, and 5 products. We choose supply $s$ and demand $d$ as random variables that follow the Pareto distribution.
Maximum Weight Matching and Stochastic Linear Program.
We explore both the maximum weight matching problem and the linear program that arises from it. Let $G=(V,E)$ be a general graph, where each edge $e\in E$ is associated with a (possibly) random weight $w_{e}$. For each node $v\in V$, denote $E(v)$ as the set of edges incident to $v$. Based on this setup, we consider the following linear program
@@
Mean-variance Portfolio Optimization.
Consider constructing a portfolio based on $m$ assets. Each asset $i$ has a rate of return $r_{i}$ that is random with mean $\mu_{i}$. The goal is to minimize the variance of the portfolio while ensuring that the expected rate of return surpasses a target level $b$. The problem can be formulated as
$\begin{array}[]{ll}\min_{\theta}&\mathbb{E}\left[(\sum_{i=1}^{m}(r_{i}-\mu_{i})\theta_{i})^{2}\right]\\[5.0pt] \text{subject to }&\sum_{i=1}^{m}\mu_{i}\theta_{i}\geq b,\\[5.0pt] &\sum_{i=1}^{m}\theta_{i}=1,\\[5.0pt] &\theta_{i}\geq 0,\quad\forall i=1,\ldots,m,\end{array}$ (68)
@@ F.3 Hyperparameter Profiling
Subsample Size.
We explored scenarios where $k$ (equivalently $k_{1}$ and $k_{2}$) is both dependent on and independent of the total sample size $n$ (see Figures 5(a), 6(a), and 6(b)). The results suggest that a constant $k$ generally suffices, although the optimal $k$ varies by problem instance. For example, Figures 6(a) and 6(b) show that $k=2$ yields the best performance; increasing $k$ degrades results. Conversely, in Figure 5(a), $k=2$ proves inadequate, with larger $k$ delivering good results. The underlying reason is that the effective performance of $\mathsf{MoVE}$ requires $\theta^{*}\in\operatorname*{arg\,max}_{\theta\in\Theta}p_{k}(\theta)$. In the former, this is achieved with only two samples, enabling $\mathsf{MoVE}$ to identify $\theta^{*}$ with a subsample size of $2$. For the latter, a higher number of samples is required to meet this condition, explaining the suboptimal performance at $k=2$. In Figure 7, we simulate $p_{k}(\theta)$ for the two cases, which further explains the influence of the subsample size.
Ensemble Size.
In Figure 8, we illustrate the performance of $\mathsf{MoVE}$ and $\mathsf{ROVE}$ under different $B,B_{1},B_{2}$, where we set $k=k_{1}=k_{2}=10$ and $\epsilon=0.005$. From the figure, we find that the performance of our ensemble methods is improving in the ensemble sizes.
Threshold $\epsilon$.
The optimal choice of $\epsilon$ in $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ is problem-dependent and related to the number of (near) optimal solutions. This dependence is illustrated by the performance of $\mathsf{ROVE}$ shown in Figures 5(b) and 5(c). Hence, we propose an adaptive strategy defined as follows: Let $g(\epsilon):=1/B_{2}\cdot\sum_{b=1}^{B_{2}}\mathbbm{1}(\hat{\theta}_{n}(\epsilon)\in\widehat{\Theta}^{\epsilon,b}_{k_{2}})$, where we use $\hat{\theta}_{n}(\epsilon)$ to emphasize the dependency of $\hat{\theta}_{n}$ on $\epsilon$. Then, we select $\epsilon^{*}:=\min\left\{\epsilon:g(\epsilon)\geq 1/2\right\}$. By definition, $g(\epsilon)$ is the proportion of times that $\hat{\theta}_{n}(\epsilon)$ is included in the “near optimum set” $\widehat{\Theta}^{\epsilon,b}_{k_{2}}$. The choice of $\epsilon^{*}$ makes it more likely for the true optimal solution to be included in the “near optimum set”, instead of being ruled out by suboptimal solutions. Practically, $\epsilon^{*}$ can be efficiently determined using a binary search as an intermediate step between Phases I and II. To prevent data leakage, we compute $\epsilon^{*}$ using $\mathbf{z}_{1:\lfloor\frac{n}{2}\rfloor}$ (Phase I data) for $\mathsf{ROVEs}$. From Figure 5, we observe that this adaptive strategy exhibits decent performance for all scenarios. Similar behaviors can also be observed for $\mathsf{ROVEs}$ in Figure 9.
Recommended Configurations.
Based on the profiling results, we summarize the recommended configurations used in all other experiments presented in the paper (unless specified otherwise):
  • For discrete space $\Theta$, use $k=\max(10,n/200),B=200$ for $\mathsf{MoVE}$, and $k_{1}=k_{2}=\max(10,n/200),B_{1}=20,B_{2}=200$ for $\mathsf{ROVE}$ and $\mathsf{ROVEs}$.
  • For continuous space $\Theta$, use $k_{1}=\max(30,n/2),k_{2}=\max(30,n/200),B_{1}=50,B_{2}=200$ for $\mathsf{ROVE}$ and $\mathsf{ROVEs}$.
  • The $\epsilon$ in $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ is selected such that $\max_{\theta\in\mathcal{S}}(1/B_{2})\sum_{b=1}^{B_{2}}\mathbbm{1}(\theta\in\widehat{\Theta}^{\epsilon,b}_{k_{2}})\approx 1/2$.
@@
F.4 Additional Experimental Results
@@
  • In Figure 18, we demonstrate how the tail heaviness of the problem affects the algorithm performance. The figure shows that the performance gap between $\mathsf{ROVE}$, $\mathsf{ROVEs}$, and the base algorithm becomes increasingly significant as the tail of the uncertainty becomes heavier. This supports the effectiveness of $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ in handling heavy-tailed uncertainty, where the base algorithm’s performance suffers. Note that here $\mathsf{MoVE}$ behaves similarly as the base learner due to optima multiplicity.
@@
Figure 17: Comparison of $\bar{\eta}_{k,\delta}$ for $\mathsf{MoVE}$ and $\mathsf{ROVE}$ in a linear program with multiple optima (corresponds to the instance in Figure 4(d)). Threshold $\epsilon$ is chosen as $\epsilon=4$ when $k=k_{1}=k_{2}=10$ and $\epsilon=2.5$ when $k=k_{1}=k_{2}=50$, according to the adaptive strategy. Note that $\bar{\eta}_{k,\delta}=\max_{\theta\in\Theta}p_{k}(\theta)-\max_{\theta\in\Theta\backslash\Theta^{\delta}}p_{k}(\theta)$ by (28), which measures the generalization sensitivity. For $\mathsf{MoVE}$, we have $p_{k}(\theta)=\mathbb{P}(\hat{\theta}_{k}^{SAA}=\theta)$; and for $\mathsf{ROVE}$, we have $p_{k}(\theta)=\mathbb{P}(\theta\in\widehat{\Theta}_{k}^{\epsilon})$, where $\widehat{\Theta}_{k}^{\epsilon}$ is the $\epsilon$-optimal set of SAA defined in (24). From the figure, we can observe that the issue brought by the presence of multiple optimal solutions can be alleviated using the two-phase strategy in $\mathsf{ROVE}$.
Figure 18: Influence of tail heaviness in the stochastic linear program with multiple optima with $n=10^{6}$. Hyperparameters: $k=50,B=2000$ for $\mathsf{MoVE}$, $k_{1}=k_{2}=50,B_{1}=200,B_{2}=5000$ for $\mathsf{ROVE}$ and $\mathsf{ROVEs}$. The tail heaviness parameter corresponds to the mean of the Pareto random coefficient.